Future Value Compute the future value in year 8 of a $350 deposit in year 3 and another $150 deposit at the end of year 5 using a 10% interest rate.

Future Value of an Annuity What is the future value of a $1,700 annuity payment over 10 years if the interest rates are 9 percent?

Present Value of a Perpetuity What's the present value, when interest rates are 6.80 percent, of a $130 payment made every year forever?

Future Value of an Annuity Due If the future value of an ordinary, 6-year annuity is $6,900 and interest rates are 7 percent, what's the future value of the same annuity due?

Future Value of Multiple Annuities Assume that you contribute $140 per month to a retirement plan for 20 years. Then you are able to increase the contribution to $240 per month for another 20 years. Given a 8.0 percent interest rate, what is the value of your retirement plan after 40 years?
Answer #1

1.
=350*1.1^5+150*1.1^3=763.3285

2.
=1700/9%*(1.09^10-1)=25827.9805200734

3.
=130/6.8%=1911.76470588235

4.
=6900*1.07=7383

5.
=140/(8%/12)*((1+8%/12)^(12*20)-1)*(1+8%/12)^(12*20)+240/(8%/12)*((1+8%/12)^(12*20)-1)=547643.13795377